A rectangular garden is 3times as long as its wide.The perimeter is 32meters.Find the dimensions of the garden.
Answer by mananth(16946)

width = x
length =3x
Perimeter of rectangle = 2(L+W)
32=2(3x+x)
32=4x
/4
x=8m
Width = 8
and length = 3*8 = 24 m